### v1

# Remove that Square Root: A New Efficient Scale-Invariant Version of AdaGrad

Abstract
Adaptive methods are extremely popular in machine learning as they make learning rate tuning less expensive. This paper introduces a novel optimization algorithm named KATE, which presents a scale-invariant adaptation of the well-known AdaGrad algorithm. We prove the scale-invariance of KATE for the case of Generalized Linear Models. Moreover, for general smooth non-convex problems, we establish a convergence rate of ${\cal O}(\nicefrac{{\log T}}{{\sqrt{T}}})$ for KATE, matching the best-known ones for AdaGrad and Adam. We also compare KATE to other state-of-the-art adaptive algorithms Adam and AdaGrad in numerical experiments with different problems, including complex machine learning tasks like image classification and text classification on real data. The results indicate that KATE consistently outperforms AdaGrad and matches/surpasses the performance of Adam in all considered scenarios.
1 Introduction
In this work, we consider the following unconstrained optimization problem:
$\min_{w\in\mathbb{R}^{d}}f(w),$ (1)
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is a $L$-smooth and generally non-convex function. In particular, we are interested in the situations when the objective has either expectation $f(w)=\mathbb{E}_{\xi\sim{\cal D}}[f_{\xi}(w)]$ or finite-sum $f(w)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(w)$ form. Such minimization problems are crucial in machine learning, where $w$ corresponds to the model parameters. Solving these problems with stochastic gradient-based optimizers has gained much interest owing to their wider applicability and low computational cost. Stochastic Gradient Descent (SGD[Robbins and Monro, 1951] and similar algorithms require the knowledge of parameters like $L$ for convergence and are very sensitive to the choice of the stepsize in general. Therefore, SGD requires hyperparameter tuning, which can be computationally expensive. To address these issues, it is common practice to use adaptive variants of stochastic gradient-based methods that can converge without knowing the function’s structure.
There exist many adaptive algorithms such as AdaGrad [Duchi et al., 2011], Adam [Kingma and Ba, 2014], AMSGrad [Reddi et al., 2019], D-Adaptation [Defazio and Mishchenko, 2023], Prodigy [Mishchenko and Defazio, 2023] and their variants. These adaptive techniques are capable of updating their step sizes on the fly. For instance, the AdaGrad method determines its step sizes using a cumulative sum of the coordinate-wise squared (stochastic) gradient of all the previous iterates:
$\text{{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}\sf{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}AdaGrad}}:}\;w_{t+1}=w_{t}-\frac{\beta g_{t}}{\sqrt{\text{ diag}\left(\Delta I+\sum\limits_{\tau=1}^{t}g_{\tau}g_{\tau}^{\top}\right)}},$ (2)
where $g_{t}$ represents an unbiased estimator of $\nabla f(w_{t})$, i.e., $\mathbb{E}\left[g_{t}\mid w_{t}\right]=\nabla f(w_{t})$, $\text{diag}(M)\in\mathbb{R}^{d}$ is a vector of diagonal elements of matrix $M\in\mathbb{R}^{d\times d}$, $\Delta>0$, and the division by vector is done component-wise. Ward et al. [2020] has shown that this method achieves a convergence rate of $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ for smooth functions, similar to SGD, without prior knowledge of the functions’ parameters. However, the performance of AdaGrad deteriorates when applied to data that may exhibit poor scaling or ill-conditioning. In this work, we propose a novel algorithm, KATE, to address the issues of poor data scaling. KATE is also a stochastic adaptive algorithm that can achieve a convergence rate of $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ for smooth non-convex functions in terms of $\min_{t\in[T]}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]^{2}$.
1.1 Related Work
A significant amount of research has been done on adaptive methods over the years, including AdaGrad [Duchi et al., 2011, McMahan and Streeter, 2010], AMSGrad [Reddi et al., 2019], RMSProp [Tieleman and Hinton, 2012], AI-SARAH [Shi et al., 2023], and Adam [Kingma and Ba, 2014]. However, all these works assume that the optimization problem is contained in a bounded set. To address this issue, Li and Orabona [2019] proposes a variant of the AdaGrad algorithm, which does not use the gradient of the last iterate (this makes the step sizes of $t$-th iteration conditionally independent of $g_{t}$) for computing the step sizes and proves convergence for the unbounded domain.
Each of these works considers a vector of step sizes for each coefficient. Duchi et al. [2011] and McMahan and Streeter [2010] simultaneously proposed the original AdaGrad algorithm. However, McMahan and Streeter [2010] was the first to consider the vanilla scalar form of AdaGrad, known as
$\text{{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}\sf{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}AdaGradNorm}}: }w_{t+1}=w_{t}-\frac{\beta g_{t}}{\sqrt{\Delta+\sum_{\tau=0}^{t}\left\|g_{\tau}\right\|^{2}}}.$ (3)
[…]
Recently, Defazio and Mishchenko [2023] introduced the D-Adaptation method, which has gathered considerable attention due to its promising empirical performances. In order to choose the adaptive step size optimally, one requires knowledge of the initial distance from the solution, i.e., $D\coloneqq\left\|w_{0}-w_{*}\right\|$ where $w_{*}\in\text{argmin}_{w\in\mathbb{R}^{d}}f(w)$. The D-Adaptation method works by maintaining an estimate of $D$ and the stepsize choice in this case is $\nicefrac{{d_{t}}}{{\sqrt{\sum_{\tau=0}^{t}\|g_{\tau}\|^{2}}}}$ for the $t$-th iteration. [^3] Here $d_{t}$ is an estimate of $D$. Mishchenko and Defazio [2023] further modifies the algorithm in a follow-up work and introduces Prodigy (with stepsize choice $\nicefrac{{d^{2}_{t}}}{{\sqrt{\sum_{\tau=0}^{t}d_{\tau}^{2}\|g_{\tau}\|^{2}}}}$) to improve the convergence speed.
Table 1: Summary of convergence guarantees for closely-related adaptive algorithms to solve smooth non-convex stochastic optimization problems. Convergence rates are given in terms of $\min_{t\in[T]}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]^{2}$. Moreover, we highlight our method’s scale-invariance property.
[htb] Algorithm Convergence rate Scale invariance AdaGradNorm [Ward et al., 2020] $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ AdaGrad [Défossez et al., 2020] $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ Adam [Défossez et al., 2020] $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ KATE (this work) $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$
Another exciting line of work on adaptive methods is Polyak stepsizes. Polyak [1969] first proposed Polyak stepsizes for subgradient methods, and recently, the stochastic version (also known as SPS) was introduced by Oberman and Prazeres [2019], Loizou et al. [2021] and Gower et al. [2021]. For a finite sum problem $\min_{w\in\mathbb{R}^{d}}f(w):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(w)$, Loizou et al. [2021] uses $\frac{f_{i}(w_{k})-f_{i}^{*}}{c\|\nabla f_{i}(w^{k})\|^{2}}$ as their stepsize choices, [^4] Here $f_{i}^{*}:=\min_{w\in\mathbb{R}^{d}}f_{i}(w)\quad\forall i\in[n]$. while Oberman and Prazeres [2019] uses $\frac{2(f(w_{k})-f^{*})}{\mathbb{E}\left[\|\nabla f_{i}(w^{k})\|^{2}\right]}$ for $k$-th iteration. However, these methods are impractical when $f^{*}$ or $f_{i}^{*}$ is unknown. Following its introduction, several variants of the SPS algorithm emerged [Li et al., 2022, D’Orazio et al., 2021]. Lately, Orvieto et al. [2022] tackled the issues with unknown $f_{i}^{*}$ and developed a truly adaptive variant. In practice, the SPS method shows excellent empirical performance on overparameterized deep learning models (which satisfy the interpolation condition i.e. $f_{i}^{*}=0\quad\forall i\in[n]$) [Loizou et al., 2021].
1.2 Main Contribution
Our main contributions are summarized below.
$\bullet$ KATE: new scale-invariant version of AdaGrad.
We propose a new method called KATE that can be seen as a version of AdaGrad, which does not use a square root in the denominator of the stepsize. To compensate for this change, we introduce a new sequence defining the numerator of the stepsize. We prove that KATE is scale-invariant for generalized linear models: if the starting point is zero, then the loss values (and training and test accuracies in the case of classification) at points generated by KATE are independent of the data scaling (Proposition 2.1), meaning that the speed of convergence of KATE is the same as for the best scaling of the data.
$\bullet$ Convergence for smooth non-convex problems.
We prove that for smooth non-convex problems with noise having bounded variance KATE has ${\cal O}(\nicefrac{{\log(T)}}{{\sqrt{T}}})$ convergence rate (Theorem 3.4), matching the best-known rates for AdaGrad and Adam [Défossez et al., 2020].
$\bullet$ Numerical experiments.
We empirically illustrate the scale-invariance of KATE on the logistic regression task and test its performance on logistic regression (see Section 4.1), image classification, and text classification problems (see Section 4.2). In all the considered scenarios, KATE outperforms AdaGrad and works either better or comparable to Adam.
1.3 Notation
[…]
2 Motivation and Algorithm Design
We focus on solving the minimization problem (1) using a variant of AdaGrad. We aim to design an algorithm that performs well, irrespective of how poorly the data is scaled.
Generalized linear models.
Here, we consider the parameter estimation problem in generalized linear models (GLMs) [Nelder and Wedderburn, 1972, Agresti, 2015] using maximum likelihood estimation. GLMs are an extension of linear models and encompass several other valuable models, such as logistic [Hosmer Jr et al., 2013] and Poisson regression [Frome, 1983], as special cases. The parameter estimation to fit GLM on dataset $\{x_{i},y_{i}\}_{i=1}^{n}$ (where $x_{i}\in\mathbb{R}^{d}$ are feature vectors and $y_{i}$ are response variables) can be reformulated as
$\min_{w\in\mathbb{R}^{d}}f(w)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\varphi_{i}\left(x_{i}^{\top}w\right)$ (4)
for differentiable functions $\varphi_{i}:\mathbb{R}\to\mathbb{R}$ [Shalev-Shwartz and Ben-David, 2014]. For example, the linear regression on data $\{x_{i},y_{i}\}_{i=1}^{n}$ is equivalent to solving (4) with $\varphi_{i}(z)=(z-y_{i})^{2}$. Next, the choice of $\varphi_{i}$ for logistic regression is $\varphi_{i}(z)=\log\left(1+\exp{\left(-y_{i}z\right)}\right)$.
Scale-invariance.
Now consider the instances of fitting GLMs on two datasets $\{x_{i},y_{i}\}_{i=1}^{n}$ and $\{Vx_{i},y_{i}\}_{i=1}^{n}$, where $V\in\mathbb{R}^{d\times d}$ is a diagonal matrix with positive entries. Note that the second dataset is a scaled version of the first one where the $k$-th component of feature vectors $x_{i}$ are multiplied by a scalar $V_{kk}$. Then, the minimization problems corresponding to datasets $\{x_{i},y_{i}\}_{i=1}^{n}$ and $\{Vx_{i},y_{i}\}_{i=1}^{n}$ are (4) and
$\min_{w\in\mathbb{R}^{d}}f^{V}(w)$$\coloneqq$ $\frac{1}{n}\sum_{i=1}^{n}\varphi_{i}\left(x_{i}^{\top}Vw\right),$ (5)
respectively, for functions $\varphi_{i}$. In this work, we want to design an algorithm with equivalent performance for the problems (4) and (5). If we can do that, the new algorithm’s performance will not deteriorate for poorly scaled data. To develop such an algorithm, we replace the denominator of AdaGrad step size with its square (remove the square root from the denominator), i.e., $\forall k\in[d]$
$w_{t+1}[k]$ $=$ $w_{t}[k]-\frac{\beta m_{t}[k]}{\sum_{\tau=0}^{t}g_{\tau}^{2}[k]}g_{t}[k].$ (6)
for some [^5] Sequence $\{m_{t}\}_{t\geq 0}$ can depend on the problem but is assumed to be scale-invariant. $m_{t}\in\mathbb{R}^{d}$. The following proposition shows that this method (6) satisfies a scale-invariance property with respect to functional value.
Proposition 2.1 (Scale invariance).
Suppose we solve problems (4) and (5) using algorithm (6). Then, the iterates $\hat{w}_{t}$ and $\hat{w}^{V}_{t}$ corresponding to (4) and (5) follow: $\forall k\in[d]$
$\hat{w}_{t+1}[k]$ $=$ $\hat{w}_{t}[k]-\frac{\beta m_{t}[k]}{\sum_{\tau=0}^{t}g_{\tau}^{2}[k]}g_{t}[k],$ (7)
$\quad\hat{w}^{V}_{t+1}[k]$ $=$ $\hat{w}^{V}_{t}[k]-\frac{\beta m_{t}[k]}{\sum_{\tau=0}^{t}\left(g_{\tau}^{V}[k]\right)^{2}}g^{V}_{t}[k]$ (8)
with $g_{\tau}=\varphi^{\prime}_{i_{\tau}}(x_{i_{\tau}}^{\top}\hat{w}_{\tau})x_{i_{\tau}}$ and $g^{V}_{\tau}=\varphi^{\prime}_{i_{\tau}}(x_{i_{\tau}}^{\top}V\hat{w}_{\tau})Vx_{i_{\tau}}$ for $i_{\tau}$ chosen uniformly from $[n]$, $\tau=0,1,\ldots,t$, $t\geq 0$. Moreover, updates (7) and (8) satisfy
$\hat{w}_{t}=V\hat{w}^{V}_{t},\quad Vg_{t}=g^{V}_{t},\quad f\left(\hat{w}_{t}\right)=f^{V}\left(\hat{w}^{V}_{t}\right)$
for all $t\geq 0$ when $\hat{w}_{0}=\hat{w}^{V}_{0}=0\in\mathbb{R}^{d}$. Furthermore we have
$\left\|g_{t}^{V}\right\|_{V^{-2}}^{2}$ $=$ $\left\|g_{t}\right\|^{2}.$ (9)
Proof.
First, we will show $\hat{w}_{t}=V\hat{w}^{V}_{t}$ and $Vg_{t}=g^{V}_{t}$ using induction. Note that for $\tau=1$ and $k\in[d]$, we get
$\hat{w}_{1}[k]$ $=$ $\frac{-\beta m_{0}[k]\varphi^{\prime}_{i_{0}}(0)x_{i_{0}}[k]}{\left(\varphi^{\prime}_{i_{0}}(0)x_{i_{0}}[k]\right)^{2}}=\frac{-\beta m_{0}[k]}{\varphi^{\prime}_{i_{0}}(0)x_{i_{0}}[k]},$
$\hat{w}^{V}_{1}[k]$ $=$ $\frac{-\beta m_{0}[k]\varphi^{\prime}_{i_{0}}(0)V_{kk}x_{i_{0}}[k]}{\left(\varphi^{\prime}_{i_{0}}(0)V_{kk}x_{i_{0}}[k]\right)^{2}}=\frac{-\beta m_{0}[k]}{\varphi^{\prime}_{i_{0}}(0)V_{kk}x_{i_{0}}[k]}.$
as $\hat{w}_{0}=\hat{w}^{V}_{0}=0$. Therefore, we have $\forall k\in[d],\hat{w}_{1}[k]=V_{kk}\hat{w}^{V}_{1}[k]$. This can be equivalently written as $\hat{w}_{1}=V\hat{w}^{V}_{1}$, as $V$ is a diagonal matrix. Then it is easy to check
$Vg_{1}$ $=$ $\varphi^{\prime}_{i_{1}}\left(x_{i_{1}}^{\top}\hat{w}_{1}\right)Vx_{i_{1}}$ (10)
$=$ $\varphi^{\prime}_{i_{1}}\left(x_{i_{1}}^{\top}V\hat{w}^{V}_{1}\right)Vx_{i_{1}}$
$=$ $g_{1}^{V},$
where the second equality follows from $\hat{w}_{1}=V\hat{w}^{V}_{1}$. Now, we assume the proposition holds for $\tau=1,\cdots,t$. Then, we need to prove this proposition for $\tau=t+1$. Note that, from (7) we have
$\hat{w}_{t+1}[k]$ $=$ $\hat{w}_{t}[k]-\frac{\beta m_{t}[k]}{\sum_{\tau=0}^{t}g^{2}_{\tau}[k]}g_{t}[k]$
$=$ $V_{kk}\hat{w}^{V}_{t}[k]-\frac{\beta m_{t}[k]V_{kk}^{2}}{\sum_{\tau=0}^{t}\left(g^{V}_{\tau}[k]\right)^{2}}\frac{g^{V}_{t}[k]}{V_{kk}}$
$=$ $V_{kk}\hat{w}^{V}_{t+1}[k].$
Here, the second last equality follows from $\hat{w}_{\tau}=V\hat{w}^{V}_{\tau}$ and $Vg_{\tau}=g^{V}_{\tau}\quad\forall\tau\in[t]$, while the last equality holds due to (8). Therefore, we have $\hat{w}_{t+1}=V\hat{w}^{V}_{t+1}$. Then similar to (10) we get $Vg_{t+1}=g^{V}_{t+1}$ using $\hat{w}_{t+1}=V\hat{w}^{V}_{t+1}$. Again, using $\hat{w}_{t}=V\hat{w}^{V}_{t}$, we can rewrite $f(\hat{w}_{t})$ as follow
$f(\hat{w}_{t})$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\varphi_{i}\left(x_{i}^{\top}\hat{w}_{t}\right)$
$=$ $\frac{1}{n}\sum_{i=1}^{n}\varphi_{i}\left(x_{i}^{\top}V\hat{w}^{V}_{t}\right)$
$=$ $f^{V}\left(\hat{w}^{V}_{t}\right).$
The last equality follows from (5). This proves $f(\hat{w}_{t})=f^{V}\left(\hat{w}^{V}_{t}\right)$. Finally using $Vg_{t}=g^{V}_{t}$ we get
$\left\|g_{t}^{V}\right\|_{V^{-2}}^{2}$ $=$ $\left(g_{t}^{V}\right)^{\top}V^{-2}g_{t}^{V}$
$=$ $g_{t}^{\top}VV^{-2}Vg_{t}$
$=$ $\left\|g_{t}\right\|^{2}.$
This completes the proof of Proposition 2.1. ∎
The Proposition 2.1 highlights that the update rule of the form (6) satisfies a scale-invariance property for GLMs.
Design of KATE.
In order to construct an algorithm following the update rule (6), one may choose $m_{t}[k]=1\;\forall k\in[d]$. However, the step size from (6) in this case may decrease very fast, and the resulting method does not necessarily converge. Therefore, we need a more aggressive choice of $m_{t}$, which grows with $t$. It motivates the construction of our algorithm KATE (Algorithm 1), [^6] Note that, for $m_{t}=b_{t}\forall t$ we get the AdaGrad algorithm. where we choose $m_{t}[k]=\sqrt{\eta[k]b_{t}^{2}[k]+\sum_{\tau=0}^{t}\frac{g_{\tau}^{2}[k]}{b_{\tau}^{2}[k]}}$. Note that the term $\sum_{\tau=0}^{t}\frac{g_{\tau}^{2}[k]}{b_{\tau}^{2}[k]}$ is scale-invariant for GLMs (follows from Proposition 2.1). To make $m_{t}$ scale-invariant, we choose $\eta\in\mathbb{R}^{d}$ in the following way:
  • $\eta\to 0$: When $\eta$ is very small, $m_{t}$ is also approximately scale-invariant for GLMs.
  • $\eta=\nicefrac{{1}}{{\left(\nabla f(w_{0})\right)^{2}}}$: In this case $\eta b_{t}^{2}=\nicefrac{{b_{t}^{2}}}{{\left(\nabla f(w_{0})\right)^{2}}}$ is scale-invariant for GLMs (follows from Proposition 2.1) as well as $m_{t}$.
Algorithm 1 KATE
0:  Initial point $w_{0}\in\mathbb{R}^{d}$, step size $\beta>0,\eta\in\mathbb{R}^{d}_{+}$ and $b_{-1},m_{-1}=0$.
1:  for $t=0,1,...,T$ do
2:     Compute $g_{t}\in\mathbb{R}^{d}$ such that $\mathbb{E}\left[g_{t}\right]=\nabla f(w_{t})$.
3:     $b_{t}^{2}=b_{t-1}^{2}+g_{t}^{2}$
4:     $m_{t}^{2}=m_{t-1}^{2}+\eta g_{t}^{2}+\frac{g_{t}^{2}}{b_{t}^{2}}$
5:     $w_{t+1}=w_{t}-\frac{\beta m_{t}}{b_{t}^{2}}g_{t}$
KATE can be rewritten in the following coordinate form
$w_{t+1}[k]=w_{t}[k]-\nu_{t}[k]g_{t}[k],\qquad\forall k\in[d],$ (11)
where $g_{t}$ is an unbiased estimator of $\nabla f(w_{t})$ and the per-coefficient step size $\nu_{t}[k]$ is defined as
$\nu_{t}[k]\coloneqq\frac{\beta\sqrt{\eta[k]b^{2}_{t}[k]+\sum_{\tau=0}^{t}\frac{g^{2}_{\tau}[k]}{b_{\tau}^{2}[k]}}}{b_{t}^{2}[k]}.$ (12)
Note that the numerator of the steps $\nu_{t}[k]$ is increasing with iterations $t$. However, one of the crucial properties of this step size choice is that the steps always decrease with $t$, which we rely on in our convergence analysis.
Lemma 2.2 (Decreasing step size).
For $\nu_{t}[k]$ defined in (11) we have
$\nu_{t+1}[k]\leq\nu_{t}[k],\qquad\forall k\in[d].$ (13)
3 Convergence Analysis
In this section, we present and discuss the convergence guarantees of KATE. In the first subsection, we list the assumptions made about the problem.
3.1 Assumptions
In all our theoretical results, we assume that $f$ is smooth as defined below.
Assumption 3.1 ($L$-smooth).
Function $f$ is $L$-smooth, i.e. for all $w,w^{\prime}\in\mathbb{R}^{d}$
$f(w^{\prime})\leq f(w)+\left\langle\nabla f(w),w^{\prime}-w\right\rangle+\frac{L}{2}\left\|w-w^{\prime}\right\|^{2}.$ (14)
This assumption is standard in the literature of adaptive methods [Li and Orabona, 2019, Ward et al., 2020, Liu et al., 2022]. Moreover, we assume that at any iteration $t$ of KATE, we can access $g_{t}$ — a noisy and unbiased estimate of $\nabla f(w_{t})$. We also make the following assumption on the noise of the gradient estimate $g_{t}$.
Assumption 3.2 (Bounded Variance).
For fixed constant $\sigma>0$, the variance of the stochastic gradient $g_{t}$ (unbiased estimate of $\left(\text{unbiased estimate of }\nabla f(w_{t})\right)$ at any time $t$ satisfies
$\mathbb{E}_{t}\left[\left\|g_{t}-\nabla f(w_{t})\right\|^{2}\right]\leq\sigma^{2}.$ (BV)
[…]
3.2 Main Results
In this section, we cover the main convergence guarantees of KATE for both deterministic and stochastic setups.
Deterministic setting.
We first present our results for the deterministic setting. In this setting, we consider the gradient estimate to have no noise (i.e. $\sigma^{2}=0$) and $g_{t}=\nabla f(w_{t})$. The main result in this setting is summarized below.
Theorem 3.3.
Suppose $f$ satisfy Assumption 3.1 and $g_{t}=\nabla f(w_{t})$. Moreover, $\beta>0$ and $\eta[k]>0$ are chosen such that $\nu_{0}[k]\leq\frac{1}{L}$ for all $k\in[d]$. Then the iterates of KATE satisfies
$\min_{t\leq T}\left\|\nabla f(w_{t})\right\|^{2}$$\leq$ $\frac{\left(\frac{2(f(w_{0})-f_{*})}{\sqrt{\eta_{0}}\beta}+\sum_{k=1}^{d}b_{0}[k]\right)^{2}}{T+1},$
where $\eta_{0}\coloneqq\min_{k\in[d]}\eta[k]$.
Discussion on Theorem 3.3.
Theorem 3.3 establishes an $\mathcal{O}\left(\nicefrac{{1}}{{T}}\right)$ convergence rate for KATE, which is optimal for finding a first-order stationary point of a non-convex problem [Carmon et al., 2020]. However, this result is not parameter-free. To prove the convergence, we assume that $\nu_{0}[k]\leq\frac{1}{L},\;\forall k\in[d]$ in Theorem 3.3, which is equivalent to $\beta\sqrt{1+\eta_{0}\left(\nabla_{k}f(w_{0})\right)^{2}}\leq\nicefrac{{\left(\nabla_{k}f(w_{0})\right)^{2}}}{{L}},\;\forall k\in[d]$. Note that the later condition holds for sufficiently small (dependent on $L$) values of $\beta,\eta_{0}>0$.
However, it is possible to derive a parameter-free version of Theorem 3.3. Indeed, Lemma 2.2 implies that the step sizes are decreasing. Therefore, we can break down the analysis of KATE into two phases: Phase I when $\nu_{0}[k]>\nicefrac{{1}}{{L}}$ and Phase II when $\nu_{0}[k]\leq\nicefrac{{1}}{{L}}$, when the current analysis works, and then follow the proof techniques of Ward et al. [2020] and Xie et al. [2020]. We leave this extension as a possible future direction of our work.
Stochastic setting.
Next, we present the convergence guarantees for KATE in the stochastic case, when we can access an unbiased gradient estimate $g_{t}$ with non-zero noise.
Theorem 3.4.
Suppose $f$ satisfy Assumption 3.1 and $g_{t}$ is an unbiased estimator of $\nabla f(w_{t})$ such that BV holds. Moreover, we assume $\|\nabla f(w_{t})\|^{2}\leq\gamma^{2}$ for all $t$. Then the iterates of KATE satisfy
$\min_{t\leq T}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]\leq\left(\frac{\|g_{0}\|}{T}+\frac{2(\gamma+\sigma)}{\sqrt{T}}\right)^{\nicefrac{{1}}{{2}}}\sqrt{\frac{2\mathcal{C}_{f}}{\beta\sqrt{\eta_{0}}}},$
where $\eta_{0}\coloneqq\min_{k\in[d]}\eta[k]$ and
$\mathcal{C}_{f}$$\coloneqq$ $f(w_{0})-f_{*}+2\beta\sigma\sum_{k=1}^{d}\sqrt{\eta[k]}\log\left(\frac{e(\sigma^{2}+\gamma^{2})T}{g_{0}^{2}[k]}\right)$
$+\sum_{k=1}^{d}\left(\frac{\beta^{2}\eta[k]L}{2}+\frac{\beta^{2}L}{2g_{0}^{2}[k]}\right)\log\left(\frac{e(\sigma^{2}+\gamma^{2})T}{g_{0}^{2}[k]}\right).$
Comparison with prior work.
Theorem 3.4 shows an $\mathcal{O}\left(\nicefrac{{\log^{\nicefrac{{1}}{{2}}}T}}{{T^{\nicefrac{{1}}{{4}}}}}\right)$ convergence rate for KATE with respect to the metric $\min_{t\leq T}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]$ for the stochastic setting. Note that, in the stochastic setting, KATE achieves a slower rate than Theorem 3.3 due to noise accumulation. Up to the logarithmic factor, this rate is optimal [Arjevani et al., 2023]. Similar rates for the same metric follow from the results [^7] Défossez et al. [2020] derive ${\cal O}(\nicefrac{{\log T}}{{\sqrt{T}}})$ convergence rates for AdaGrad and Adam in terms of $\min_{t\leq T}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|^{2}\right]$ which is not smaller than $\min_{t\leq T}\left(\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]\right)^{2}$. of [Défossez et al., 2020] for AdaGrad and Adam.
Finally, Li and Orabona [2019] considers a variant of AdaGrad closely related to KATE:
$w_{t+1}=w_{t}-\frac{\beta g_{t}}{\left(\text{ diag}\left(\Delta I+\sum_{\tau=1}^{t-1}g_{\tau}g_{\tau}^{\top}\right)\right)^{\frac{1}{2}+\varepsilon}},$ (15)
for some $\varepsilon\in[0,\nicefrac{{1}}{{2}})$ and $\Delta>0$. It differs from AdaGrad in two key aspects: the denominator of the stepsize does not contain the last stochastic gradient, and also, instead of the square root of the sum of squared gradients, this sum is taken in the power of $\nicefrac{{1}}{{2}}+\varepsilon$. However, the results from Li and Orabona [2019] do not imply convergence for the case of $\varepsilon$, which is expected since, in this case, the stepsize converges to zero too quickly in general. To compensate for such a rapid decrease, in KATE, we introduce an increasing sequence $m_{t}$ in the numerator of the stepsize.
Proof technique.
Compared to the AdaGrad, KATE uses more aggressive steps (the larger numerator of KATE due to the extra term $\sum_{\tau=0}^{t}\nicefrac{{g^{2}_{\tau}[k]}}{{b^{2}_{\tau}[k]}}$). Therefore, we expect KATE to have better empirical performance. However, introducing $\sum_{\tau=0}^{t}\nicefrac{{g^{2}_{\tau}[k]}}{{b^{2}_{\tau}[k]}}$ in the numerator raises additional technical difficulties in the proof technique. Fortunately, as we rigorously show, the KATE steps $\nu_{t}[k]$ retain some of the critical properties of AdaGrad steps. For instance, they (i) are lower bounded by AdaGrad steps up to a constant, (ii) decrease with iteration $t$ (Lemma 2.2), and (iii) have closed-form upper bounds for $\sum_{t=0}^{T}\nu^{2}_{t}[k]g^{2}_{t}[k]$. These are indeed the primary building blocks of our proof technique.
(a) Plot of $f(w_{t})$
(b) Plot of Accuracy
(c) Plot of Gradient Norm
Figure 1: Comparison of KATE on scaled and un-scaled data. Figures 0(a), and 0(b) plot the functional value $f(w_{t})$ and accuracy on scaled and unscaled data, respectively. Figure 0(c) plots $\left\|\nabla f(w_{t})\right\|^{2}$ and $\left\|\nabla f(w_{t})\right\|_{V^{-2}}^{2}$ for unscaled and scaled data respectively.
4 Numerical Experiments
In this section, we implement KATE in several machine learning tasks to evaluate its performance. The code to reproduce our results can be found at https://github.com/nazya/KATE.
4.1 Logistic Regression
In this section, we consider the logistic regression model
$\min_{w\in\mathbb{R}^{d}}f(w)=\frac{1}{n}\sum_{i=1}^{n}\log\left(1+\exp\left(-y_{i}x_{i}^{\top}w\right)\right),$ (16)
to elaborate on the scale-invariance and robustness of KATE for various initializations.
Synthetic data.
To conduct this experiment, we set the total number of samples to $1000$ (i.e. $n=1000$). Here, we simulate the independent vectors $x_{i}\in\mathbb{R}^{20}$ such that each entry is from $\mathcal{N}(0,1)$. Moreover, we generate a diagonal matrix $V\in\mathbb{R}^{20\times 20}$ such that $\log V_{kk}\overset{\mathrm{iid}}{\sim}\text{Unif}(-10,10),\ \forall k\in[20]$. Similarly, we generate $w^{*}\in\mathbb{R}^{20}$ with each component from $\mathcal{N}(0,1)$ and set the labels
$y_{i}=\begin{cases}1&x_{i}^{\top}Vw^{*}\geq 0,\\ -1&x_{i}^{\top}Vw^{*}<0\end{cases}\qquad\forall i\in[n].$
4.1.1 Scale Invariance
In this experiment, we implement ${\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}\sf{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}KATE}}$ on problems (4) (for unscaled data) and (5) (for scaled data) with $\varphi_{i}(z)=\log\left(1+\exp{\left(-y_{i}z\right)}\right)$. We run ${\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}\sf{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}KATE}}$ for $10,000$ iterations with mini-batch size $10$, $\eta=\nicefrac{{1}}{{\left(\nabla f(w_{0})\right)^{2}}}$ and plot functional value $f(w_{t})$ and accuracy in Figures 0(a) and 0(b). We use the proportion of correctly classified samples to compute accuracy, i.e. $\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\left\{y_{i}x_{i}^{\top}w_{t}\geq 0\right\}}$. In plots 0(a) and 0(b), the functional value and accuracy of ${\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}\sf{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}KATE}}$ coincide, which aligns with our theoretical findings (Proposition 2.1). Figure 0(c) plots $\left\|\nabla f(w_{t})\right\|^{2}$ and $\left\|\nabla f(w_{t})\right\|_{V^{-2}}^{2}$ for unscaled and scaled data respectively. Here, (9) explains the identical values taken by the corresponding gradient norms of KATE iterates for the scaled and unscaled data.
4.1.2 Robustness of KATE
We compare KATE’s performance with four other algorithms: AdaGrad, AdaGradNorm, SGD-decay and SGD-constant, similar to the section 5.1 of Ward et al. [2020]. For each algorithm, we initialize with $w_{0}=0\in\mathbb{R}^{20}$ and independently draw a sample of mini-batch size $10$ to update the weight vector $w_{t}$. We compare the algorithms $\bullet$ AdaGrad with stepsize $\frac{\beta}{\sqrt{\Delta+\sum_{\tau=0}^{t}g^{2}_{\tau}}}$, $\bullet$ AdaGradNorm with step size $\frac{\beta}{\sqrt{\Delta+\sum_{\tau=0}^{t}\left\|g_{\tau}\right\|^{2}}}$, $\bullet$ SGD-decay with stepsize $\nicefrac{{\beta}}{{\Delta\sqrt{t+1}}}$, and $\bullet$ SGD-constant with step size $\nicefrac{{\beta}}{{\Delta}}$. Similarly, for KATE we use stepsize $\frac{\beta m_{t}}{b_{t}^{2}}$ where $m_{t}^{2}=\eta b_{t}^{2}+\sum_{\tau=0}^{t}\nicefrac{{g_{\tau}^{2}}}{{b_{\tau}^{2}}}$ and $b_{t}^{2}=\Delta+\sum_{\tau=0}^{t}g_{\tau}^{2}$. Here, we choose $\beta=f(w_{0})-f(w^{*})$ and vary $\Delta$ in $\{10^{-8},10^{-6},10^{-4},10^{-2},1,10^{2},10^{4},10^{6},10^{8}\}$.
In Figures 1(a), 1(b), and 1(c), we plot the functional value $f(w_{t})$ (on the $y$-axis) after $10^{4},5\times 10^{4}$, and $10^{5}$ iterations, respectively. In theory, the convergence of SGD requires the knowledge of smoothness constant $L$. Therefore, when the $\Delta$ is small (hence the stepsize is large), SGD-decay and SGD-constant diverge. However, the adaptive algorithms KATE, AdaGrad, and AdaGradNorm can auto-tune themselves and converge for a wide range of $\Delta$s (even when the $\Delta$ is too small). As we observe in Figure 2, when the $\Delta$ is small, KATE outperforms all other algorithms. For instance, when $\Delta=10^{-8}$, KATE achieves a functional value of $10^{-3}$ after only $10^{4}$ iterations (see Figure 1(a)), while other algorithms fail to achieve this even after $10^{5}$ iterations (see Figure 1(c)). Furthermore, KATE performs as well as AdaGrad and better than other algorithms when the $\Delta$ is large. In particular, this experiment highlights that KATE is robust to initialization $\Delta$.
(a) Plot of $f(w_{t})$
(b) Plot of Accuracy
(c) Plot of Gradient Norm
Figure 2: Comparison of KATE with AdaGrad, AdaGradNorm, SGD-decay and SGD-constant for different values of $\Delta$ (on $x$-axis for logistic regression model. Figure 1(a), 1(b) and 1(c) plots the functional value $f(w_{t})$ (on $y$-axis) after $10^{4},5\times 10^{4}$, and $10^{5}$ iterations respectively.
(a) Dataset: heart
(b) Dataset: australian
(c) Dataset: splice
(d) Dataset: heart
(e) Dataset: australian
(f) Dataset: splice
Figure 3: Comparison of KATE with AdaGrad, AdaGradNorm, SGD-decay and SGD-constant on datasets heart, australian, and splice from LIBSVM. Figures 2(a), 2(b) and 2(c) plot the functional value $f(w_{t})$, while 2(d), 2(e) and 2(f) plot the accuracy on $y$-axis for $5,000$ iterations.
4.1.3 Peformance of KATE on Real Data
In this section, we examine KATE’s performance on real data. We test KATE on three datasets: heart, australian, and splice from the LIBSVM library. The response variables $y_{i}$ of each of these datasets contain two classes, and we use them for binary classification tasks using a logistic regression model (16). We take $\eta=\nicefrac{{1}}{{\left(\nabla f(w_{0})\right)^{2}}}$ for KATE and tune $\beta$ in all the experiments. For tuning $\beta$, we do a grid search on the list $\{10^{-10},10^{-8},10^{-6},10^{-4},10^{-2},1\}$. Similarly, we tune stepsizes for other algorithms. We take $5$ trials for each of these algorithms and plot the mean of their trajectories.
We plot the functional value $f(w_{t})$ (i.e. loss function) in Figures 2(a), 2(b) and 2(c), whereas Figures 2(d), 2(e) and 2(f) plot the corresponding accuracy of the weight vector $w_{t}$ on the $y$-axis for $5,000$ iterations. We observe that KATE performs superior to all other algorithms, even on real datasets.
Figure 4: CIFAR10: $\eta=0$
Figure 5: CIFAR10: $\eta=0.001$
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: CIFAR10: $\eta=0$
Figure 5: CIFAR10: $\eta=0.001$
Figure 6: CIFAR10: $\eta=0.1$
Figure 7: Emotion: $\eta=0$
Figure 8: Emotion: $\eta=0.001$
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 7: Emotion: $\eta=0$
Figure 8: Emotion: $\eta=0.001$
Figure 9: Emotion: $\eta=0.1$
4.2 Training of Neural Networks
In this section, we compare the performance of KATE, AdaGrad and Adam on two tasks, i.e. training ResNet18 [He et al., 2016] on the CIFAR10 dataset [Krizhevsky and Hinton, 2009] and BERT [Devlin et al., 2018] fine-tuning on the emotions dataset [Saravia et al., 2018] from the Hugging Face Hub. We choose standard parameters for Adam ($\beta_{1}=0.9$ and $\beta_{2}=0.999$) that are default values in PyTorch and select the learning rate of $10^{-5}$ for all considered methods. We run KATE with different values of $\eta\in\{0,10^{-1},10^{-2}\}$. For the image classification task, we normalize the images (similar to Horváth and Richtárik [2020]) and use a mini-batch size of 500. For the BERT fine-tuning, we use a mini-batch size 160 for all methods.
Figures 6-9 report the evolution of top-1 accuracy and cross-entropy loss (on the $y$-axis) calculated on the test data. For the image classification task, we observe that KATE with different choices of $\eta$ outperforms Adam and AdaGrad. Finally, we also observe that KATE performs comparably to Adam on the BERT fine-tuning task and is better than AdaGrad. These preliminary results highlight the potential of KATE to be applied for training neural networks for different tasks.
Supplementary Material
We organize the Supplementary Material as follows: Section A presents some technical lemmas required for our analysis. In section B, we provide the proofs for the main results of our work.
Appendix A Technical Lemmas
Lemma A.1 (AM-GM).
For $\lambda>0$ we have
$ab\leq\frac{\lambda}{2}a^{2}+\frac{1}{2\lambda}b^{2}.$ (17)
Lemma A.2 (Cauchy-Schwarz Inequality).
For $a_{1},\cdots,a_{n},b_{1},\cdots,b_{n}\in\mathbb{R}$ we have
$\left(\sum_{i=1}^{n}a_{i}^{2}\right)\left(\sum_{i=1}^{n}b_{i}^{2}\right)$$\geq$ $\left(\sum_{i=1}^{n}a_{i}b_{i}\right)^{2}.$ (18)
Lemma A.3 (Holder’s Inequality).
Suppose $X,Y$ are two random variables and $p,q>1$ satisfy $\frac{1}{p}+\frac{1}{q}=1$. Then
$\mathbb{E}\left(|XY|\right)\leq\left(\mathbb{E}\left(|X|^{p}\right)\right)^{\frac{1}{p}}\left(\mathbb{E}\left(|Y|^{q}\right)\right)^{\frac{1}{q}}.$ (19)
Lemma A.4 (Jensen’s Inequality).
For a convex function $g:\mathbb{R}^{d}\to\mathbb{R}$ and a random variable $X$ such that $\mathbb{E}{\left(\Psi(X)\right)}$ and $\Psi\left(\mathbb{E}{\left(X\right)}\right)$ are finite, we have
$\Psi\left(\mathbb{E}{\left(X\right)}\right)\leq\mathbb{E}{\left(\Psi(X)\right)}.$ (20)
Lemma A.5.
For $a_{1},a_{2},\cdots,a_{n}\geq 0$ and $b_{1},b_{2},\cdots,b_{n}>0$ we have
$\sum_{i=1}^{n}\frac{a_{i}}{\sqrt{b_{i}}}\geq\frac{\sum_{i=1}^{n}a_{i}}{\sqrt{\sum_{i=1}^{n}b_{i}}}.$ (21)
Proof.
Expanding the LHS of (21) we get
$\left(\sum_{i=1}^{n}\frac{a_{i}}{\sqrt{b_{i}}}\right)^{2}$ $=$ $\sum_{i=1}^{n}\frac{a_{i}^{2}}{b_{i}}+2\sum_{i\neq j}\frac{a_{i}a_{j}}{\sqrt{b_{i}b_{j}}}$ (22)
$\geq$ $\sum_{i=1}^{n}\frac{a_{i}^{2}}{b_{i}}.$
The last inequality follows from $\frac{a_{i}}{\sqrt{b_{i}}}\geq 0$ for all $i\in[n]$. Now, using Cauchy-Schwarz Inequality (18), we have
$\left(\sum_{i=1}^{n}\frac{a_{i}^{2}}{b_{i}}\right)\left(\sum_{i=1}^{n}b_{i}\right)$$\geq$ $\left(\sum_{i=1}^{n}a_{i}\right)^{2}.$ (23)
Then combining (22) and (23), we get
$\left(\sum_{i=1}^{n}\frac{a_{i}}{\sqrt{b_{i}}}\right)^{2}\left(\sum_{i=1}^{n}b_{i}\right)$$\geq$ $\left(\sum_{i=1}^{n}a_{i}\right)^{2}.$
Finally dividing both sides by $\sum_{i=1}^{n}b_{i}$ and taking square root we get the desired result. ∎
Lemma A.6.
For $k\in[d]$ and $t\geq 1$ we have
$\mathbb{E}_{t}\left[\left(\frac{\beta\sqrt{\eta[k]}}{\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}-\nu_{t}[k]\right)\nabla_{k}f(w_{t})g_{t}[k]\right]$$\leq$ $\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{2\sqrt{b^{2}_{t-1}[k]+\left(\nabla f(w_{t})\right)^{2}+\sigma^{2}}}$ (24)
$+2\beta\sqrt{\eta[k]}\sigma\mathbb{E}_{t}\left[\frac{g^{2}_{t}[k]}{b_{t}^{2}[k]}\right]$
Proof.
Note that, using $\nu_{t}[k]\geq\frac{\beta\sqrt{\eta[k]}}{b_{t}[k]}$ we have
$\frac{\beta\sqrt{\eta[k]}}{\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}-\nu_{t}[k]$ (25)
$\leq$ $\beta\sqrt{\eta[k]}\left(\frac{1}{\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}-\frac{1}{b_{t}[k]}\right)$
$=$ $\beta\sqrt{\eta[k]}\left(\frac{b_{t}^{2}[k]-b_{t-1}^{2}[k]-(\nabla_{k}f(w_{t}))^{2}-\sigma^{2}}{b_{t}[k]\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}\left(b_{t}[k]+\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}\right)}\right)$
$=$ $\beta\sqrt{\eta[k]}\left(\frac{g_{t}^{2}[k]-(\nabla_{k}f(w_{t}))^{2}-\sigma^{2}}{b_{t}[k]\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}\left(b_{t}[k]+\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}\right)}\right)$
$=$ $\beta\sqrt{\eta[k]}\left(\frac{\left(g_{t}[k]+\nabla_{k}f(w_{t})\right)\left(g_{t}[k]-\nabla_{k}f(w_{t})\right)-\sigma^{2}}{b_{t}[k]\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}\left(b_{t}[k]+\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}\right)}\right)$
$\leq$ $\frac{\beta\sqrt{\eta[k]}\left|\left(g_{t}[k]+\nabla_{k}f(w_{t})\right)\left(g_{t}[k]-\nabla_{k}f(w_{t})\right)\right|}{b_{t}[k]\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}\left(b_{t}[k]+\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}\right)}$
$+\frac{\beta\sqrt{\eta[k]}\sigma^{2}}{b_{t}[k]\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}\left(b_{t}[k]+\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}\right)}$
$\leq$ $\frac{\beta\sqrt{\eta[k]}\left|g_{t}[k]-\nabla_{k}f(w_{t})\right|}{b_{t}[k]\sqrt{b^{2}_{t-1}[k]+(\nabla_{k}f(w_{t}))^{2}+\sigma^{2}}}+\frac{\beta\sqrt{\eta[k]}\sigma}{b_{t}[k]\sqrt{b^{2}_{t-1}[k]+(\nabla_{k}f(w_{t}))^{2}+\sigma^{2}}}.$
Note that the second last inequality follows from the use of triangle inequality in the following way
$\left(g_{t}[k]+\nabla_{k}f(w_{t})\right)\left(g_{t}[k]-\nabla_{k}f(w_{t})\right)-\sigma^{2}$$\leq$ $\left|\left(g_{t}[k]+\nabla_{k}f(w_{t})\right)\left(g_{t}[k]-\nabla_{k}f(w_{t})\right)-\sigma^{2}\right|$
$\leq$ $\left|\left(g_{t}[k]+\nabla_{k}f(w_{t})\right)\left(g_{t}[k]-\nabla_{k}f(w_{t})\right)\right|+\sigma^{2},$
while the last inequality follows from
$b_{t}[k]+\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}$$\geq$ $\left|g_{t}[k]\right|+\left|\nabla_{k}f(w_{t})\right|\geq\left|g_{t}[k]+\nabla_{k}f(w_{t})\right|,$
$b_{t}[k]+\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}$$\geq$ $\sigma.$
Then from (25) we have
$\mathbb{E}_{t}\left[\left(\frac{\beta\sqrt{\eta[k]}}{\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}-\nu_{t}[k]\right)\nabla_{k}f(w_{t})g_{t}[k]\right]$ (26)
$\leq$ $\underbrace{\beta\sqrt{\eta[k]}\mathbb{E}_{t}\left[\frac{\left|g_{t}[k]-\nabla_{k}f(w_{t})\right|\left|\nabla_{k}f(w_{t})\right|\left|g_{t}[k]\right|}{b_{t}[k]\sqrt{b^{2}_{t-1}[k]+(\nabla_{k}f(w_{t}))^{2}+\sigma^{2}}}\right]}_{\textrm{term I}}$
$+\underbrace{\beta\sqrt{\eta[k]}\mathbb{E}_{t}\left[\frac{\sigma\left|\nabla_{k}f(w_{t})\right|\left|g_{t}[k]\right|}{b_{t}[k]\sqrt{b^{2}_{t-1}[k]+(\nabla_{k}f(w_{t}))^{2}+\sigma^{2}}}\right]}_{\textrm{term II}}.$
For term I in (26), we use Lemma A.1 with
$\lambda$ $=$ $\frac{2\sigma^{2}}{\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}},$
$a$ $=$ $\frac{\left|g_{t}[k]\right|}{b_{t}[k]},$
$b$ $=$ $\frac{\left|g_{t}[k]-\nabla_{k}f(w_{t})\right|\left|\nabla_{k}f(w_{t})\right|}{\sqrt{b^{2}_{t-1}[k]+(\nabla_{k}f(w_{t}))^{2}+\sigma^{2}}},$
to get
$\beta\sqrt{\eta[k]}\mathbb{E}_{t}\left[\frac{\left|g_{t}[k]-\nabla_{k}f(w_{t})\right|\left|\nabla_{k}f(w_{t})\right|\left|g_{t}[k]\right|}{b_{t}[k]\sqrt{b^{2}_{t-1}[k]+(\nabla_{k}f(w_{t}))^{2}+\sigma^{2}}}\right]$ (27)
$\leq$ $\frac{\beta\sqrt{\eta[k]}\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}{4\sigma^{2}}\frac{\left(\nabla_{k}f(w_{t})\right)^{2}\mathbb{E}_{t}\left[g_{t}[k]-\nabla_{k}f(w_{t})\right]^{2}}{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}$
$+\frac{\beta\sqrt{\eta[k]}\sigma^{2}}{\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}\mathbb{E}_{t}\left[\frac{g^{2}_{t}[k]}{b^{2}_{t}[k]}\right]$
$\leq$ $\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{4\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}+\beta\sqrt{\eta[k]}\sigma\mathbb{E}_{t}\left[\frac{g^{2}_{t}[k]}{b_{t}^{2}[k]}\right].$
The last inequality follows from BV. Similarly, we again use Lemma A.1 with
$\lambda$ $=$ $\frac{2}{\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}},$
$a$ $=$ $\frac{\sigma\left|g_{t}[k]\right|}{b_{t}[k]},$
$b$ $=$ $\frac{\left|\nabla_{k}f(w_{t})\right|}{\sqrt{b^{2}_{t}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}$
and $\sqrt{b^{2}_{t}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}\geq\sigma$ to get
$\beta\sqrt{\eta[k]}\mathbb{E}_{t}\left[\frac{\sigma\left|\nabla_{k}f(w_{t})\right|\left|g_{t}[k]\right|}{b_{t}[k]\sqrt{b^{2}_{t-1}[k]+(\nabla_{k}f(w_{t}))^{2}+\sigma^{2}}}\right]\leq\beta\sqrt{\eta[k]}\sigma\mathbb{E}_{t}\left[\frac{g^{2}_{t}[k]}{b_{t}^{2}[k]}\right]+\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{4\sqrt{b^{2}_{t-1}[k]+\left(\nabla f(w_{t})\right)^{2}+\sigma^{2}}}.$ (28)
Therefore using (27) and (28) in (27) we get
$\mathbb{E}_{t}\left[\left(\frac{\beta\sqrt{\eta[k]}}{\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}-\nu_{t}[k]\right)\nabla_{k}f(w_{t})g_{t}[k]\right]$$\leq$ $2\beta\sqrt{\eta[k]}\sigma\mathbb{E}_{t}\left[\frac{g^{2}_{t}[k]}{b_{t}^{2}[k]}\right]$
$+\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{2\sqrt{b^{2}_{t-1}[k]+\left(\nabla f(w_{t})\right)^{2}+\sigma^{2}}}.$
This completes the proof of this Lemma. ∎
Lemma A.7.
$\sum_{t=0}^{T}\frac{g_{t}^{2}[k]}{b_{t}^{2}[k]}\leq\log\left(\frac{b_{T}^{2}[k]}{b_{0}^{2}[k]}\right)+1$ (29)
Proof.
Using $b_{t}^{2}[k]=\sum_{\tau=0}^{t}g_{\tau}^{2}[k]$ we have
$\sum_{t=0}^{T}\frac{g_{t}^{2}[k]}{b_{t}^{2}[k]}$ $=$ $1+\sum_{t=1}^{T}\frac{g_{t}^{2}[k]}{b_{t}^{2}[k]}$
$=$ $1+\sum_{t=1}^{T}\frac{b_{t}^{2}[k]-b_{t-1}^{2}[k]}{b_{t}^{2}[k]}$
[…]
$\leq$ $1+\sum_{t=1}^{T}\int_{b^{2}_{t-1}[k]}^{b_{t}^{2}[k]}\frac{dz}{z}$
$=$ $1+\int_{b^{2}_{0}[k]}^{b_{T}^{2}[k]}\frac{dz}{z}$
$=$ $1+\log\left(\frac{b_{T}^{2}[k]}{b_{0}^{2}[k]}\right).$
The inequality follows from the fact $\frac{1}{b_{t}^{2}[k]}\leq\frac{1}{z}$ when $b_{t-1}^{2}[k]\leq z\leq b_{t}^{2}[k]$. This completes the proof of the Lemma. ∎
Appendix B Proof of Main Results
B.1 Proof of Lemma 2.2
Lemma B.1 (Decreasing step size).
For $\nu_{t}[k]$ defined in (11) we have
$\nu_{t+1}[k]\leq\nu_{t}[k]\qquad\forall k\in[d].$
Proof.
We want to show that $\nu_{t+1}[k]\leq\nu_{t}[k]$. Taking square and rearranging the terms (13) is equivalent to proving
$b_{t}^{4}[k]m_{t+1}^{2}[k]\leq b_{t+1}^{4}[k]m_{t}^{2}[k].$ (30)
Using the expansion of $m^{2}_{t+1}[k],b^{2}_{t+1}[k]$, LHS of (30) can be expanded as follow
$b_{t}^{4}[k]m_{t+1}^{2}[k]$ $=$ $b_{t}^{4}[k]\left(m_{t}^{2}[k]+\eta[k]g^{2}_{t+1}[k]+\frac{g_{t+1}^{2}[k]}{b_{t}^{2}[k]+g^{2}_{t+1}[k]}\right).$ (31)
Similarly, the RHS of (30) can be expanded to
$b_{t+1}^{4}[k]m_{t}^{2}[k]$ $=$ $m_{t}^{2}[k]\left(b_{t}^{2}[k]+g_{t+1}^{2}[k]\right)^{2}$ (32)
$=$ $m_{t}^{2}[k]b_{t}^{4}[k]+m^{2}_{t}[k]g_{t+1}^{4}[k]+2m_{t}^{2}[k]g_{t+1}^{2}[k]b_{t}^{2}[k].$
Therefore using (31) and (32), inequality (30) is equivalent to
$b_{t}^{4}[k]\left(m_{t}^{2}[k]+\eta[k]g^{2}_{t+1}[k]+\frac{g_{t+1}^{2}[k]}{b_{t}^{2}[k]+g^{2}_{t+1}[k]}\right)$$\leq$ $m_{t}^{2}[k]b_{t}^{4}[k]+m^{2}_{t}[k]g_{t+1}^{4}[k]+2m_{t}^{2}[k]g_{t+1}^{2}[k]b_{t}^{2}[k].$ (33)
Now subtracting $m_{t}^{2}[k]b_{t}^{4}[k]$ from both sides of (33) and then multiplying both sides by $b_{t}^{2}[k]+g^{2}_{t+1}[k]$, (33) is equivalent to
$\eta[k]g_{t+1}^{2}[k]b_{t}^{6}[k]+\eta[k]g_{t+1}^{4}[k]b_{t}^{4}[k]+g_{t+1}^{2}[k]b_{t}^{4}[k]$$\leq$ $m_{t}^{2}[k]g_{t+1}^{4}[k]b_{t}^{2}[k]+2m_{t}^{2}[k]g_{t+1}^{2}[k]b^{4}_{t}[k]+m_{t}^{2}[k]g^{6}_{t+1}[k]$ (34)
$+2m_{t}^{2}[k]g_{t+1}^{4}[k]b_{t}^{2}[k].$
Therefore, proving (13) is equivalent to proving (34). Note that, from the expansion $m_{t}^{2}[k]=\eta[k]b_{t}^{2}[k]+\sum_{\tau=0}^{t}\frac{g_{t}^{2}[k]}{b_{t}^{2}[k]}$, we have $m_{t}^{2}[k]\geq\frac{g_{0}^{2}[k]}{b_{0}^{2}[k]}=1$ and $m_{t}^{2}[k]\geq\eta[k]b_{t}^{2}[k]$. Then using $m_{t}^{2}[k]\geq 1$ we get
$g_{t+1}^{4}[k]b_{t}^{2}[k]\leq m_{t}^{2}[k]g_{t+1}^{4}[k]b_{t}^{2}[k].$ (35)
Again, using $m_{t}^{2}[k]\geq\eta[k]b_{t}^{2}[k]$, we have
$\eta[k]g_{t+1}^{2}[k]b_{t}^{6}[k]+\eta[k]g_{t+1}^{4}[k]b_{t}^{4}[k]$$\leq$ $m_{t}^{2}[k]g_{t+1}^{2}[k]b^{4}_{t}[k]+m_{t}^{2}[k]g_{t+1}^{4}[k]b_{t}^{2}[k].$ (36)
Then adding (35) and (36) we get
$\eta[k]g_{t+1}^{2}[k]b_{t}^{6}[k]+\eta[k]g_{t+1}^{4}[k]b_{t}^{4}[k]+g_{t+1}^{2}[k]b_{t}^{4}[k]$$\leq$ $m_{t}^{2}[k]g_{t+1}^{4}[k]b_{t}^{2}[k]+2m_{t}^{2}[k]g_{t+1}^{2}[k]b^{4}_{t}[k].$ (37)
Therefore, (34) is true due to (37) and $m_{t}^{2}[k]g^{6}_{t+1}[k]+2m_{t}^{2}[k]g_{t+1}^{4}[k]b_{t}^{2}[k]\geq 0$. This completes the proof of the Lemma. ∎
B.2 Proof of Theorem 3.3
Theorem B.2.
Suppose $f$ is $L$-smooth, $g_{t}=\nabla f(w_{t})$ and $\eta,\beta$ are chosen such that $\nu_{0}[k]\leq\frac{1}{L}$ for all $k\in[d]$. Then for (11) we have
$\min_{t\leq T}\left\|\nabla f(w_{t})\right\|^{2}\leq\frac{1}{T+1}\left(\sum_{k=1}^{d}b_{0}[k]+\frac{2(f(w_{0})-f_{*})}{\sqrt{\eta}\beta}\right)^{2}.$
Proof.
Suppose $g_{t}=\nabla f(w_{t})$. Then using the smoothness of $f$ we get
$f(w_{T+1})$$\leq$ $f(w_{T})+\left\langle g_{T},w_{T+1}-w_{T}\right\rangle+\frac{L}{2}\left\|w_{T+1}-w_{T}\right\|^{2}$
$=$ $f(w_{T})+\sum_{k=1}^{d}g_{T}[k]\left(w_{T+1}[k]-w_{T}[k]\right)+\frac{L}{2}\sum_{k=1}^{d}\left(w_{T+1}[k]-w_{T}[k]\right)^{2}$
$=$ $f(w_{T})-\sum_{k=1}^{d}\nu_{T}[k]g_{T}^{2}[k]+\frac{L}{2}\sum_{k=1}^{d}\nu_{T}^{2}[k]g_{T}^{2}[k]$
$=$ $f(w_{T})-\sum_{k=1}^{d}\nu_{T}[k]\left(1-\nu_{T}[k]\frac{L}{2}\right)g_{T}^{2}[k].$
Then using this bound recursively we get
$f(w_{T+1})$$\leq$ $f(w_{0})-\sum_{t=0}^{T}\sum_{k=1}^{d}\nu_{t}[k]\left(1-\nu_{t}[k]\frac{L}{2}\right)g_{t}^{2}[k].$
Note that, we initialized KATE such that $\nu_{0}[k]\leq\frac{1}{L}\forall k\in[d]$. Therefore using Lemma 2.2 we have $\nu_{t}[k]\leq\frac{1}{L}$, which is equivalent to $1-\nu_{t}[k]\frac{L}{2}\geq\frac{1}{2}$ for all $k\in[d]$. Hence from (B.2) we have
$f(w_{T+1})$$\leq$ $f(w_{0})-\sum_{t=0}^{T}\sum_{k=1}^{d}\frac{\nu_{t}[k]}{2}g_{t}^{2}[k].$
Then rearranging the terms and using $f(w_{T+1})\geq f_{*}$ we get
$\sum_{t=0}^{T}\sum_{k=1}^{d}\frac{\nu_{t}[k]}{2}g_{t}^{2}[k]\leq f(w_{0})-f_{*}.$ (38)
Then from (38) and $m_{t}[k]\geq\sqrt{\eta_{0}}b_{t}[k]$ we get
$\sum_{t=0}^{T}\sum_{k=1}^{d}\frac{g_{t}^{2}[k]}{b_{t}[k]}\leq\frac{2(f(w_{0})-f_{*})}{\sqrt{\eta_{0}}\beta}.$ (39)
Now from the definition of $b_{t}^{2}[k]$, we have $b^{2}_{t}[k]=b_{t-1}^{2}[k]+g_{t}^{2}[k]$. This can be rearranged to get
$b_{T}[k]$ $=$ $b_{T-1}[k]+\frac{g_{T}^{2}[k]}{b_{T}[k]+b_{T-1}[k]}$ (40)
$\leq$ $b_{T-1}[k]+\frac{g_{T}^{2}[k]}{b_{T}[k]}$
$\leq$ $b_{0}[k]+\sum_{t=0}^{T}\frac{g_{t}^{2}[k]}{b_{t}[k]}.$ (41)
Here the last inequality (41) follows from recursive use of (40). Then, taking squares on both sides and summing over $k\in[d]$ we get
$\sum_{k=1}^{d}b_{T}^{2}[k]$$\leq$ $\sum_{k=1}^{d}\left(b_{0}[k]+\sum_{t=0}^{T}\frac{g_{t}^{2}[k]}{b_{t}[k]}\right)^{2}$ (42)
$\leq$ $\left(\sum_{k=1}^{d}b_{0}[k]+\sum_{t=0}^{T}\sum_{k=1}^{d}\frac{g_{t}^{2}[k]}{b_{t}[k]}\right)^{2}$
$\leq$ $\left(\sum_{k=1}^{d}b_{0}[k]+\frac{2(f(w_{0})-f_{*})}{\sqrt{\eta_{0}}\beta}\right)^{2}.$
The second inequality follows from $b_{0}[k]+\sum_{t=0}^{T}\frac{g_{t}^{2}[k]}{b_{t}[k]}\geq 0$ for all $k\in[d]$ and the last inequality from (39). Now note that $\sum_{t=0}^{T}\left\|g_{t}\right\|^{2}=\sum_{t=0}^{T}\sum_{k=1}^{d}g_{t}^{2}[k]=\sum_{k=1}^{d}b^{2}_{t}[k]$. Therefore dividing both sides of (42) by $T+1$, we get
$\min_{t\leq T}\left\|\nabla f(w_{t})\right\|^{2}\leq\frac{1}{T+1}\left(\sum_{k=1}^{d}b_{0}[k]+\frac{2(f(w_{0})-f_{*})}{\sqrt{\eta_{0}}\beta}\right)^{2}.$
This completes the proof of the theorem. ∎
B.3 Proof of Theorem 3.4
Theorem B.3.
Suppose $f$ is a $L$-smooth function and $g_{t}$ is an unbiased estimator of $\nabla f(w_{t})$ such that BV holds. Moreover, we assume $\|\nabla f(w_{t})\|^{2}\leq\gamma^{2}$ for all $t$. Then KATE satisfies
$\min_{t\leq T}\mathbb{E}{\left\|\nabla f(w_{t})\right\|}$$\leq$ $\left(\frac{\|g_{0}\|}{T}+\frac{2(\gamma+\sigma)}{\sqrt{T}}\right)^{\nicefrac{{1}}{{2}}}\sqrt{\frac{2\mathcal{C}_{f}}{\beta\sqrt{\eta_{0}}}}$
where
$\mathcal{C}_{f}=f(w_{0})-f_{*}+\sum_{k=1}^{d}\left(2\beta\sqrt{\eta[k]}\sigma+\frac{\beta^{2}\eta[k]L}{2}+\frac{\beta^{2}L}{2g_{0}^{2}[k]}\right)\left(\log\left(\frac{(\sigma^{2}+\gamma^{2})T}{g_{0}^{2}[k]}\right)+1\right).$
Proof.
Using smoothness, we have
$f(w_{t+1})$$\leq$ $f(w_{t})+\left\langle\nabla f(w_{t}),w_{t+1}-w_{t}\right\rangle+\frac{L}{2}\left\|w_{t+1}-w_{t}\right\|^{2}$
$=$ $f(w_{t})+\sum_{k=1}^{d}\nabla_{k}f(w_{t})\left(w_{t+1}[k]-w_{t}[k]\right)+\frac{L}{2}\sum_{k=1}^{d}\left(w_{t+1}[k]-w_{t}[k]\right)^{2}$
$=$ $f(w_{t})-\sum_{k=1}^{d}\nu_{t}[k]\nabla_{k}f(w_{t})g_{t}[k]+\frac{L}{2}\sum_{k=1}^{d}\nu_{t}^{2}[k]g_{t}^{2}[k].$
Then, taking the expectation conditioned on $w_{t}$, we have
$\mathbb{E}_{t}\left[f(w_{t+1})\right]$$\leq$ $f(w_{t})-\sum_{k=1}^{d}\mathbb{E}_{t}\left[\nu_{t}[k]\nabla_{k}f(w_{t})g_{t}[k]\right]+\frac{L}{2}\sum_{k=1}^{d}\mathbb{E}_{t}\left[\nu_{t}^{2}[k]g_{t}^{2}[k]\right]$
$=$ $f(w_{t})-\sum_{k=1}^{d}\mathbb{E}_{t}\left[\nu_{t}[k]\nabla_{k}f(w_{t})g_{t}[k]\right]+\frac{L}{2}\sum_{k=1}^{d}\mathbb{E}_{t}\left[\nu_{t}^{2}[k]g_{t}^{2}[k]\right]$
$-\sum_{k=1}^{d}\frac{\beta\sqrt{\eta[k]}}{\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}\mathbb{E}_{t}\left[\nabla_{k}f(w_{t})\left(\nabla_{k}f(w_{t})-g_{t}[k]\right)\right]$
$=$ $f(w_{t})+\sum_{k=1}^{d}\mathbb{E}_{t}\left[\left(\frac{\beta\sqrt{\eta[k]}}{\sqrt{b_{t-1}^{2}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}-\nu_{t}[k]\right)\nabla_{k}f(w_{t})g_{t}[k]\right]$
$+\frac{L}{2}\sum_{k=1}^{d}\mathbb{E}_{t}\left[\nu_{t}^{2}[k]g_{t}^{2}[k]\right]-\sum_{k=1}^{d}\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}.$
The second last equality follows from $\mathbb{E}_{t}\left[\nabla_{k}f(w_{t})\left(\nabla_{k}f(w_{t})-g_{t}[k]\right)\right]=\nabla_{k}f(w_{t})\left(\nabla_{k}f(w_{t})-\mathbb{E}_{t}\left[g_{t}[k]\right]\right)=0$. Now we use (24) to get
$\mathbb{E}_{t}\left[f(w_{t+1})\right]$$\leq$ $f(w_{t})+\sum_{k=1}^{d}2\beta\sqrt{\eta[k]}\sigma\mathbb{E}_{t}\left[\frac{g^{2}_{t}[k]}{b_{t}^{2}[k]}\right]+\frac{L}{2}\sum_{k=1}^{d}\mathbb{E}_{t}\left[\nu_{t}^{2}[k]g_{t}^{2}[k]\right]-\sum_{k=1}^{d}\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{2\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}.$
Then rearranging the terms we have
$\sum_{k=1}^{d}\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{2\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}$$\leq$ $f(w_{t})-\mathbb{E}_{t}\left[f(w_{t+1})\right]+\sum_{k=1}^{d}2\beta\sqrt{\eta[k]}\sigma\mathbb{E}_{t}\left[\frac{g^{2}_{t}[k]}{b_{t}^{2}[k]}\right]+\frac{L}{2}\sum_{k=1}^{d}\mathbb{E}_{t}\left[\nu_{t}^{2}[k]g_{t}^{2}[k]\right].$
Now we take the total expectations to derive
$\sum_{k=1}^{d}\mathbb{E}\left[\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{2\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}\right]$$\leq$ $\mathbb{E}\left[f(w_{t})\right]-\mathbb{E}\left[f(w_{t+1})\right]+\sum_{k=1}^{d}2\beta\sqrt{\eta[k]}\sigma\mathbb{E}\left[\frac{g^{2}_{t}[k]}{b_{t}^{2}[k]}\right]+\frac{L}{2}\sum_{k=1}^{d}\mathbb{E}\left[\nu_{t}^{2}[k]g_{t}^{2}[k]\right].$
The above inequality holds for any $t$. Therefore summing up from $t=0$ to $t=T$ and using $f(w_{T+1})\geq f_{*}$ we get
$\sum_{t=0}^{T}\sum_{k=1}^{d}\mathbb{E}\left[\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{2\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}\right]$$\leq$ $f(w_{0})-f_{*}+\sum_{t=0}^{T}\sum_{k=1}^{d}2\beta\sqrt{\eta[k]}\sigma\mathbb{E}\left[\frac{g^{2}_{t}[k]}{b_{t}^{2}[k]}\right]$ (43)
$+\frac{L}{2}\sum_{t=0}^{T}\sum_{k=1}^{d}\mathbb{E}\left[\nu_{t}^{2}[k]g_{t}^{2}[k]\right].$
Note that, using the expansion of $\nu_{t}^{2}[k]$ we have
$\nu_{t}^{2}[k]$ $=$ $\frac{\beta^{2}\eta[k]b_{t}^{2}[k]+\beta^{2}\sum_{j=0}^{t}\frac{g_{j}^{2}[k]}{b_{j}^{2}[k]}}{b_{t}^{4}[k]}$ (44)
$=$ $\frac{\beta^{2}\eta[k]}{b_{t}^{2}[k]}+\frac{\beta^{2}}{b_{t}^{4}[k]}\sum_{j=0}^{t}\frac{g_{j}^{2}[k]}{b_{j}^{2}[k]}$
$\leq$ $\frac{\beta^{2}\eta[k]}{b_{t}^{2}[k]}+\frac{\beta^{2}}{b_{t}^{4}[k]b_{0}^{2}[k]}\sum_{j=0}^{t}g_{j}^{2}[k]$
$=$ $\frac{\beta^{2}\eta[k]}{b_{t}^{2}[k]}+\frac{\beta^{2}}{b_{t}^{2}[k]g_{0}^{2}[k]}.$ (45)
Here (44) follows from $b_{j}^{2}[k]\geq b_{0}^{2}[k]$ and (45) from $b_{t}^{2}[k]=\sum_{j=0}^{t}g_{j}^{2}[k]$. Then using (45) in (43) we derive
$\sum_{t=0}^{T}\sum_{k=1}^{d}\mathbb{E}\left[\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{2\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}\right]$$\leq$ $f(w_{0})-f_{*}+\sum_{t=0}^{T}\sum_{k=1}^{d}\left(2\beta\sqrt{\eta[k]}\sigma+\frac{\beta^{2}\eta[k]L}{2}+\frac{\beta^{2}L}{2g_{0}^{2}[k]}\right)\mathbb{E}\left[\frac{g^{2}_{t}[k]}{b_{t}^{2}[k]}\right]$
$\leq$ $f(w_{0})-f_{*}$
$+\sum_{k=1}^{d}\left(2\beta\sqrt{\eta[k]}\sigma+\frac{\beta^{2}\eta[k]L}{2}+\frac{\beta^{2}L}{2g_{0}^{2}[k]}\right)\mathbb{E}\left[\log\left(\frac{b^{2}_{T}[k]}{b_{0}^{2}[k]}\right)+1\right].$
Here the last inequality follows from (29). Now using Jensen’s Inequality (20) with $\Psi(z)=\log(z)$ we have
$\sum_{t=0}^{T}\sum_{k=1}^{d}\mathbb{E}\left[\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{2\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}\right]$$\leq$ $f(w_{0})-f_{*}$
$+\sum_{k=1}^{d}\left(2\beta\sqrt{\eta[k]}\sigma+\frac{\beta^{2}\eta[k]L}{2}+\frac{\beta^{2}L}{2g_{0}^{2}[k]}\right)\left(\log\left(\frac{\mathbb{E}\left[b^{2}_{T}[k]\right]}{b_{0}^{2}[k]}\right)+1\right).$
Now note that $\mathbb{E}\left[b_{T}^{2}[k]\right]=\sum_{t=0}^{T}\mathbb{E}\left[g_{t}^{2}[k]\right]=\sum_{t=0}^{T}\mathbb{E}\left[g_{t}[k]-\nabla_{k}f(w_{t})\right]^{2}+\left(\nabla_{k}f(w_{t})\right)^{2}\leq(\sigma^{2}+\gamma^{2})T$. Therefore, we have the bound
$\sum_{t=0}^{T}\sum_{k=1}^{d}\mathbb{E}\left[\frac{\beta\sqrt{\eta[k]}\left(\nabla_{k}f(w_{t})\right)^{2}}{2\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}\right]$$\leq$ $f(w_{0})-f_{*}+2\beta\sigma\sum_{k=1}^{d}\sqrt{\eta[k]}\log\left(\frac{e(\sigma^{2}+\gamma^{2})T}{b_{0}^{2}[k]}\right)$ (46)
$+\sum_{k=1}^{d}\left(\frac{\beta^{2}\eta[k]L}{2}+\frac{\beta^{2}L}{2g_{0}^{2}[k]}\right)\log\left(\frac{e(\sigma^{2}+\gamma^{2})T}{b_{0}^{2}[k]}\right).$
Here the RHS is exactly $\mathcal{C}_{f}$. Using (21) we have
$\sum_{k=1}^{d}\frac{\left(\nabla_{k}f(w_{t})\right)^{2}}{\sqrt{b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}$$\geq$ $\frac{\sum_{k=1}^{d}\left(\nabla_{k}f(w_{t})\right)^{2}}{\sqrt{\sum_{k=1}^{d}b^{2}_{t-1}[k]+\left(\nabla_{k}f(w_{t})\right)^{2}+\sigma^{2}}}$ (47)
$=$ $\frac{\left\|\nabla f(w_{t})\right\|^{2}}{\sqrt{\left\|b_{t-1}\right\|^{2}+\left\|\nabla f(w_{t})\right\|^{2}+d\sigma^{2}}}.$
Therefore using (47) in (46) we arrive at
$\sum_{t=0}^{T}\mathbb{E}\left[\frac{\left\|\nabla f(w_{t})\right\|^{2}}{\sqrt{\left\|b_{t-1}\right\|^{2}+\left\|\nabla f(w_{t})\right\|^{2}+d\sigma^{2}}}\right]$$\leq$ $\frac{2\mathcal{C}_{f}}{\beta\sqrt{\eta_{0}}}.$ (48)
Now we use Holder’s Inequality (19) $\frac{\mathbb{E}(XY)}{\left(\mathbb{E}{\left|Y\right|^{3}}\right)^{\frac{1}{3}}}\leq\left(\mathbb{E}|X|^{\frac{3}{2}}\right)^{\frac{2}{3}}$ with
$X=\left(\frac{\left\|\nabla f(w_{t})\right\|^{2}}{\sqrt{\left\|b_{t-1}\right\|^{2}+\left\|\nabla f(w_{t})\right\|^{2}+d\sigma^{2}}}\right)^{\frac{2}{3}}\quad\text{ and }\quad Y=\left(\sqrt{\left\|b_{t-1}\right\|^{2}+\left\|\nabla f(w_{t})\right\|^{2}+d\sigma^{2}}\right)^{\frac{2}{3}}$
to get a lower bound on LHS of (48):
$\mathbb{E}\left[\frac{\left\|\nabla f(w_{t})\right\|^{2}}{\sqrt{\left\|b_{t-1}\right\|^{2}+\left\|\nabla f(w_{t})\right\|^{2}+d\sigma^{2}}}\right]$$\geq$ $\frac{\mathbb{E}\left[\|\nabla f(w_{t})\|^{\frac{4}{3}}\right]^{\frac{3}{2}}}{\sqrt{\mathbb{E}\left(\left\|b_{t-1}\right\|^{2}+\left\|\nabla f(w_{t})\right\|^{2}+d\sigma^{2}\right)}}$ (49)
$\geq$ $\frac{\mathbb{E}\left[\|\nabla f(w_{t})\|^{\frac{4}{3}}\right]^{\frac{3}{2}}}{\sqrt{\left\|b_{0}\right\|^{2}+2t(\gamma^{2}+d\sigma^{2})}}.$
Therefore from (48) and (49) we get
$\frac{T}{\sqrt{\|b_{0}\|^{2}+2T(\gamma^{2}+d\sigma^{2})}}\min_{t\leq T}\mathbb{E}\left[\|\nabla f(w_{t})\|^{\frac{4}{3}}\right]^{\frac{3}{2}}$$\leq$ $\frac{2\mathcal{C}_{f}}{\beta\sqrt{\eta_{0}}}.$
Then multiplying both sides by $\frac{\|b_{0}\|+\sqrt{2T}(\gamma+\sqrt{d}\sigma)}{T}$ we have
$\min_{t\leq T}\mathbb{E}\left[\|\nabla f(w_{t})\|^{\frac{4}{3}}\right]^{\frac{3}{2}}$$\leq$ $\left(\frac{\|b_{0}\|}{T}+\frac{2(\gamma+\sigma)}{\sqrt{T}}\right)\frac{2\mathcal{C}_{f}}{\beta\sqrt{\eta_{0}}}.$
Here we use $\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]^{\frac{4}{3}}\leq\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|^{\frac{4}{3}}\right]$ (follows from Jensen’s Inequality (20) with $\Psi(z)=z^{\nicefrac{{4}}{{3}}}$) in the above equation to get
$\min_{t\leq T}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]^{2}$$\leq$ $\left(\frac{\|b_{0}\|}{T}+\frac{2(\gamma+\sigma)}{\sqrt{T}}\right)\frac{2\mathcal{C}_{f}}{\beta\sqrt{\eta_{0}}}.$
This completes the proof of the Theorem. ∎